lim n→∞ (2 n+1 + 3 n+1 )/(2 n + 3 n ) is equal to (a) 2 (b)3 (c) 5 (D) none. explain pls.

lim n→∞ (2n+1 + 3n+1)/(2n + 3n) is equal to (a) 2 (b)3 (c) 5 (D) none. explain pls.

Grade:12th Pass

3 Answers

RAJORSHI PAUL
11 years ago

2^(n+1)+3^(n+1) =3^(n+1){(2/3)^(n+1) +1}

2^n+3^n =3^n{(2/3)^n +1}

As n tends to infinity (2/3)^n tends to 0

So the required limit is 3

 

Chetan Mandayam Nayakar
11 years ago

(2n+1 +3n+1)/(2n +3n)=2n+1/(2n + 3n)+3n+1/(2n + 3n)=2/(1+(3/2)n)+3/(3+(2/3)n)=0+1=1

 

jitender lakhanpal
11 years ago

Dear Aditi,

 

correct answer is 3 as (2/3)n   where n is ∞ is 0

so if we take 3n+1  from numerator and 3n  is denominator so we get (2/3)n  form

 

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

jitender